graph the circle.. x^2 + y^2 - 4x + 8y = 24
Answer by toidayma(44)

You have to change the equation to the form below:
(x-a)^2 + (y-b)^2 = R^2, whereas: O(a,b) is the coordinate of the center of the circle, and the R is the radius.
x^2 + y^2 - 4x + 8y = 24 <->
x^2 -4x +4 + y^2 + 8y + 16 = 24 + 4 +16 <->
(x - 2)^2 + (y+ 4)^2 = 44
Therefore, the center of the circle is (2,-4), and the radius is root(44).
Now you can graph it easily, right?